OpenStudy (anonymous):

WILL FAN AND MEDAL Simplify 8u ------- 10(u+v) I think it's 4 ----- 5v

OpenStudy (anonymous):

@SeanStewart16

OpenStudy (anonymous):

imposible, 4/5 is good, but u/(u+v)!!! NUUU jajja \[\frac{ a+b }{ C } = \frac{ a }{ C } + \frac{ b }{ C }\] but \[\frac{ a }{ b+c } = nothing\] maybe \[\frac{ a }{ b+c } = \frac{ 1 }{ \frac{ b+c }{ a } }=\frac{ 1 }{ \frac{ b }{ a }+\frac{ c }{ a } }\]

OpenStudy (anonymous):

Wait so.. I'm not sure what that would make the variables with 4 --- 5

OpenStudy (anonymous):

4/5 is fixed and is ok the rest you can't do nothing. \[\frac{4}{5 } ·\frac{ u }{ u+v }\] obviusly i don't fully understand the problem

OpenStudy (anonymous):

*obviously

OpenStudy (anonymous):

*didn't understood ... i'm not a good english speaker :P ar all

OpenStudy (anonymous):

It's ok :) @AihberKhan @KendrickLamar2014 @james1769 do you know what to do next?

OpenStudy (anonymous):

do you have the full problem?

OpenStudy (anonymous):

This is the full problem

OpenStudy (anonymous):

ahhh ok then it is solved! check with WolphramAlpha if you want more confidence

OpenStudy (aihberkhan):

Okay you have to first simply: u/10

OpenStudy (aihberkhan):

So the equation after this step should be: \[(8 * \frac{ u }{ 10 }) * ( u + v) \]

OpenStudy (aihberkhan):

Now after solving that by distributing, order of operations, etc, you should get: \[\frac{ 4u * u + v }{ 5 }\]

OpenStudy (anonymous):

noooooooooo that is imposible dear aihber!

OpenStudy (anonymous):

you want a common factor of 8u/10 then the other term should be 1/(u+v) \[8u/10 · \frac{1}{(u+v)}\]

OpenStudy (anonymous):

@AihberKhan it was incorrect..

OpenStudy (anonymous):

is not fair, i'm a college student soo desproportional, don't deserve any medal xD your friend made a good try

OpenStudy (anonymous):

byee